Calculus II help please!!

A tank is half full of a liquid that has a density of 900 kg/m3. Find the work W required to pump the liquid out of the spout. (Use 9.8 m/s2 for g. Assume r = 15 m andh = 5 m.)
Please show using integration. Thank you
Answer #1

1: computing the volume of the spherical tank = 4/3?r³ = V
2a. compute the total mass of the liquid in spherical tank = 900V
2b. compute the total weight of liquid in spherical tank = 900Vg
3. compute the work done by lifting this liquid's Center of Mass thru a distance = r+h
ANS = Work = 900Vg(r+h)

answered by: Litha